# 3D Scene Prompting for Scene-Consistent Camera-Controllable Video Generation

### v1

@@ 4.1 Implementation Details
We build upon CogVideoX-I2V-5B (Yang et al., 2024), extending its single-image conditioning to accept dual spatio-temporal inputs with minimal architectural changes. The key modification is repurposing the existing image conditioning channel to accept concatenated latents from both temporal frames and spatial projections. Specifically, we provide the last $w=9$ frames from $\mathbf{V}_{\text{in}}$ as temporal conditioning and $T$ projected views from the static point cloud as spatial conditioning. This enables the DiT backbone to remain entirely unchanged, preserving all pretrained video priors. Both conditions are encoded through the frozen 3D VAE and concatenated channel-wise such that $\mathbf{Z}_{\text{cond}}=\text{Concat}[\mathcal{E}(\text{Temporal}(w)),\mathcal{E}(\text{Spatial}(T))]$.
@@
We fully fine-tune the model for a total of 4K iterations with a batch size of 8 using 4 H100 GPUs, which required approximately 48 hours. We used the 16-bit Adam optimizer with a learning rate of $1\times 10^{-5}$, and adopted the same hyperparameter settings as those used in the training of CogVideoX (Yang et al., 2024). For the temporal sliding window, we provide the last 9 frames of the input video, setting $w=9$. For the projection of top-$n$ spatially adjacent views, we set $n=7$.
@@
We evaluate our method across four key aspects: camera controllability, video quality, scene consistency, and geometric consistency. Since no prior work directly addresses scene-consistent camera-controllable video generation, we compare against two categories of baselines: (1) camera-controllable methods (CameraCtrl (He et al., 2024), MotionCtrl (Wang et al., 2024b), FloVD (Jin et al., 2025), AC3D (Bahmani et al., 2025)) for camera control and video quality metrics, and (2) DFoT (Song et al., 2025), which attempts scene-consistent camera-controllable generation, for spatial and geometric consistency metrics.
We primarily evaluate on 1,000 dynamic videos from DynPose-100K (Rockwell et al., 2025). For scene consistency evaluation, we additionally test on 1,000 static videos from RealEstate10K (Zhou et al., 2018), as static scenes provide clearer spatial consistency assessment.
@@
Table 1: Evaluation of spatial and geometric consistency. We compare DFoT and Ours on the RealEstate10K (Zhou et al., 2018) and DynPose-100K (Rockwell et al., 2025) datasets. For spatial consistency, we evaluate PSNR, SSIM, and LPIPS on revisited camera trajectories, while for geometric consistency, we report the MEt3R (Asim et al., 2025) metric.
@@ 4.2 Scene-consistent video generation
As mentioned in Section 3.1, one of the unique and key challenges in scene-consistent camera-controllable video generation is maintaining spatial consistency over extended durations. From a given input video, we evaluate spatial consistency by generating camera trajectories that revisit the viewpoints in the given video. By matching frames in the generated video and the input video that share the same viewpoint, we calculate PSNR, SSIM, and LPIPS. For RealEstate10K, we evaluate the whole image, whereas we only evaluate the static regions by masking out the dynamic regions for DynPose-100K. We also assess geometric consistency using Met3R (Asim et al., 2025), which measures multi-view alignment of generated frames under the recovered camera pose.
@@ Results.
Figure 6: Visualization of generated videos following trajectories that revisit early frames in the input video. We visualize and compare frames obtained from DFoT (Song et al., 2025) and Ours. We condition both DFoT and ours to generate a video that follows the camera trajectory shown in the top row (GT Future Video), which revisits a viewpoint in the input (Input Video). The comparison shows that ours shows much more consistent generation, whereas DFoT fails to generate scene-consistent frames mainly due to the limited number of frames it can condition on. Also, the comparison of the camera trajectory shows that our method more faithfully follows the given camera condition.
Figure 7: Visualization of scene-consistent camera-controllable video generation. Comparison of different methods  (Wang et al., 2024b; He et al., 2024; Jin et al., 2025; Bahmani et al., 2025) for generating videos from the same input (shown in Input Video) that follow the camera trajectory shown in the rightmost column (Camera). Our method best preserves scene consistency with the input video. Note the red-boxed regions: while the input video shows a white wall, competing methods either lose scene detail or fail to maintain the original scene structure. In contrast, our approach accurately remembers the white wall and maintains consistent scene elements throughout generation. In addition, when compared with the GT Future Video, ours best follows the camera condition, effectively verifying the strength of our framework for scene-consistent camera-controllable video generation.
@@ 4.3 Camera-Controllable Video Generation
We employ the evaluation protocol of previous methods (He et al., 2024; Zheng et al., 2024; Jin et al., 2025) for the camera controllability. We provide an input image along with associated camera parameters for I2V models (He et al., 2024; Wang et al., 2024b; Jin et al., 2025) and solely provide camera parameters for the T2V model (Bahmani et al., 2025). For our model, we provide the last 9 frames of the input video together with the camera parameters. To evaluate how faithfully the generated video follows the camera condition, we estimate camera parameters from the synthesized video using MegaSAM (Li et al., 2024), and compare the estimated camera parameters against the condition camera trajectory $\mathbf{C}$.
As done in previous works (Hong et al., 2024b; a; Rockwell et al., 2025), the comparison between the estimated and input camera parameters is quantified using three metrics: mean rotation error (mRotErr), mean translation error (mTransErr), and mean error in the camera extrinsic matrices (mCamMC).
@@
MotionCtrl Wang et al. (2024b) 3.5654 7.8231 9.7834
CameraCtrl He et al. (2024) 3.3273 9.5989 11.2122
FloVD Jin et al. (2025) 3.4811 11.0302 12.6202
AC3D Bahmani et al. (2025) 3.0675 9.7044 11.1634
DFoT Song et al. (2025) 2.3977 8.0866 9.2330
3DScenePrompt (Ours) 2.3772 7.4174 8.6352
For the generated video, we also assess video synthesis performance using the Fréchet Video Distance (FVD) (Skorokhodov et al., 2022) and seven metrics from VBench++ (Huang et al., 2024): subject consistency, background consistency, aesthetic quality, imaging quality, temporal flickering, motion smoothness, and dynamic degree.
@@ Results.
We first evaluate camera controllability and compare our method with competitive baselines. As shown in Tab. 2, our approach consistently outperforms existing methods, indicating 3DScenePrompt is capable of generating videos with precise camera control. We then assess the overall video quality (Tab. 3) and provide qualitative comparisons (Fig. 7). As observed in Tab. 3, our method achieves the best generation quality across all metrics for dynamic video generation, which is further supported by the visual results in Fig. 7.
Table 3: Evaluation of video generation quality. We assess the quality of generated videos using FVD and VBench++ scores. For FVD, lower values indicate higher video quality. For VBench++ scores, higher values indicate better performance for all metrics. All VBench++ scores are normalized.
@@
MotionCtrl Wang et al. (2024b) 1017.4247 0.5625 0.5158 0.7093 0.3157 0.3149 0.8297 0.8432 0.7900
CameraCtrl He et al. (2024) 737.0506 0.6280 0.6775 0.8238 0.3736 0.3888 0.6837 0.6955 0.9900
FloVD Jin et al. (2025) 171.2697 0.7273 0.7964 0.8457 0.4722 0.5546 0.7842 0.8364 0.9900
AC3D Bahmani et al. (2025) 281.2140 0.7428 0.8360 0.8674 0.4766 0.5381 0.8020 0.8673 1.0000

### v2

@@ 4.1 Implementation Details
We build upon CogVideoX-I2V-5B (yang2024cogvideox), extending its single-image conditioning to accept dual spatio-temporal inputs with minimal architectural changes. The key modification is repurposing the existing image conditioning channel to accept concatenated latents from both temporal frames and spatial projections. Specifically, we provide the last $w=9$ frames from $\mathbf{V}_{\text{in}}$ as temporal conditioning and $T$ projected views from the static point cloud as spatial conditioning. This enables the DiT backbone to remain entirely unchanged, preserving all pretrained video priors. Both conditions are encoded through the frozen 3D VAE and concatenated channel-wise such that $\mathbf{Z}_{\text{cond}}=\mathcal{E}[\text{Concat}(\text{Temporal}(w),\text{Spatial}(T))]$.
@@
We fully fine-tune the model for a total of 4K iterations with a batch size of 8 using 4 H100 GPUs, which required approximately 48 hours. We used the 16-bit Adam optimizer with a learning rate of $1\times 10^{-5}$, and adopted the same hyperparameter settings as those used in the training of CogVideoX (yang2024cogvideox). For the temporal sliding window, we provide the last 9 frames of the input video, setting $w=9$. For the projection of top-$n$ spatially adjacent views, we set $n=7$.
@@
We evaluate our method across four key aspects: camera controllability, video quality, scene consistency, and geometric consistency. Since no prior work directly addresses scene-consistent camera-controllable video generation, we compare against two categories of baselines: (1) camera-controllable methods (CameraCtrl (he2024cameractrl), MotionCtrl (wang2024motionctrl), FloVD (jin2025flovd), AC3D (Bahmani_2025_CVPR)) for camera control and video quality metrics, and (2) DFoT (song2025history), which attempts scene-consistent camera-controllable generation, for spatial and geometric consistency metrics.
We primarily evaluate on 1,000 dynamic videos from DynPose-100K (rockwell2025dynamic). For scene consistency evaluation, we additionally test on 1,000 static videos from RealEstate10K (zhou2018stereo), as static scenes provide clearer spatial consistency assessment.
@@
Table 1: Evaluation of spatial and geometric consistency. We compare DFoT and our framework on the RealEstate10K (zhou2018stereo) and DynPose-100K (rockwell2025dynamic) datasets. For spatial consistency, we evaluate PSNR, SSIM, and LPIPS on revisited camera trajectories, while for geometric consistency, we report the MEt3R (asim2025met3r) metric.
@@ 4.2 Scene-consistent video generation
As mentioned in Section 3.1, one of the unique and key challenges in scene-consistent camera-controllable video generation is maintaining spatial consistency over extended durations. From a given input video, we evaluate spatial consistency by generating camera trajectories that revisit the viewpoints in the given video. By matching frames in the generated video and the input video that share the same viewpoint, we calculate PSNR, SSIM, and LPIPS. For RealEstate10K, we evaluate the whole image, whereas we only evaluate the static regions by masking out the dynamic regions for DynPose-100K. We also assess geometric consistency using Met3R (asim2025met3r), which measures multi-view alignment of generated frames under the recovered camera pose.
@@ Results.
Figure 6: Visualization of generated videos following trajectories that revisit early frames in the input video. We visualize and compare frames obtained from DFoT (song2025history) and Ours. We condition both DFoT and ours to generate a video that follows the camera trajectory shown in the top row (GT Future Video), which revisits a viewpoint in the input (Input Video). The comparison shows that ours shows much more consistent generation, whereas DFoT fails to generate scene-consistent frames mainly due to the limited number of frames it can condition on. Also, the comparison of the camera trajectory shows that our method more faithfully follows the given camera condition.
Figure 7: Visualization of scene-consistent camera-controllable video generation. Comparison of different methods  (wang2024motionctrl; he2024cameractrl; jin2025flovd; bahmani2025ac3d) for generating videos from the same input (shown in Input Video) that follow the camera trajectory shown in the rightmost column (Camera). Our method best preserves scene consistency with the input video. Note the red-boxed regions: while the input video shows a white wall, competing methods either lose scene detail or fail to maintain the original scene structure. In contrast, our approach accurately remembers the white wall and maintains consistent scene elements throughout generation. In addition, when compared with the GT Future Video, ours best follows the camera condition, effectively verifying the strength of our framework for scene-consistent camera-controllable video generation.
@@ 4.3 Camera-Controllable Video Generation
MotionCtrl wang2024motionctrl 3.5654 7.8231 9.7834
CameraCtrl he2024cameractrl 3.3273 9.5989 11.2122
FloVD jin2025flovd 3.4811 11.0302 12.6202
AC3D Bahmani_2025_CVPR 3.0675 9.7044 11.1634
DFoT song2025history 2.3977 8.0866 9.2330
3DScenePrompt ($w=1$) 2.3898 7.7819 8.9785
3DScenePrompt ($w=9$) 2.3772 7.4174 8.6352
We employ the evaluation protocol of previous methods (he2024cameractrl; zheng2024cami2v; jin2025flovd) for the camera controllability. We provide an input image along with associated camera parameters for I2V models (he2024cameractrl; wang2024motionctrl; jin2025flovd) and solely provide camera parameters for the T2V model (Bahmani_2025_CVPR). We conduct experiments using two variants of our framework, leveraging different numbers of frames ($w$) for temporal conditioning baselines, $w=1$ and $w=9$. The $w=1$ uses only the last single frame as temporal conditioning, whereas the $w=9$ model takes the last nine frames as temporal conditioning. To evaluate how faithfully the generated video follows the camera condition, we estimate camera parameters from the synthesized video using MegaSAM (li2024megasam), and compare the estimated camera parameters against the condition camera trajectory $\mathbf{C}$.
The comparison between the estimated and input camera parameters is quantified using three metrics: mean rotation error (mRotErr), mean translation error (mTransErr), and mean error in the camera extrinsic matrices (mCamMC). For the generated video, we also assess video synthesis performance using the Fréchet Video Distance (FVD)(skorokhodov2022stylegan) and seven metrics from VBench++(huang2024vbench++): subject consistency, background consistency, aesthetic quality, imaging quality, temporal flickering, motion smoothness, and dynamic degree.
@@ Results.
We first evaluate camera controllability and compare our method with competitive baselines. As shown in Tab. 2, our approach consistently outperforms existing methods, indicating 3DScenePrompt is capable of generating videos with precise camera control. We also note that the effect of using different temporal conditioning window sizes is minimal for the camera-controllability performance, suggesting that better camera-controllability is achieved from our spatial prompts rather than the increased temporal context. We then assess the overall video quality (Tab. 3) and provide qualitative comparisons (Fig. 7). As observed in Tab. 3, our method achieves the best generation quality across all metrics for dynamic video generation, which is further supported by the visual results in Fig. 7.
Table 3: Evaluation of video generation quality. We assess the quality of generated videos using FVD and VBench++ scores. For FVD, lower values indicate higher video quality. For VBench++ scores, higher values indicate better performance. All VBench++ scores are normalized.
@@
MotionCtrl (wang2024motionctrl) 1017.4247 0.5625 0.5158 0.7093 0.3157 0.3149 0.8297 0.8432 0.7900
CameraCtrl (he2024cameractrl) 737.0506 0.6280 0.6775 0.8238 0.3736 0.3888 0.6837 0.6955 0.9900
FloVD (jin2025flovd) 171.2697 0.7273 0.7964 0.8457 0.4722 0.5546 0.7842 0.8364 0.9900
AC3D (Bahmani_2025_CVPR) 281.2140 0.7428 0.8360 0.8674 0.4766 0.5381 0.8020 0.8673 1.0000
@@ 5 Conclusion
Appendix
Notation Summary
Symbol Meaning
$V_{\text{in}}\in\mathbb{R}^{L\times H\times W\times 3}$ input arbitrary-length video with $L$ frames.
$V_{\text{out}}\in\mathbb{R}^{T\times H\times W\times 3}$ generated future video with $T$ frames.
$\mathcal{T}$ text prompt for video generation for video models based on text-to-video (T2V) generation.
$\mathbf{C}=\{C_{t}\}_{t=1}^{T}$ desired camera trajectory the generated video $V_{\text{out}}$ should follow.
$C_{t}$ camera extrinsics parameter where $C_{t}\in\mathbb{SE}(3)$.
$K$ camera intrinsics parameter.
$\mathcal{F}(\cdot)$ camera-controllable video generation framework.
$\mathcal{G}(\cdot)$ video-to-future-video generation framework.
$\mathbf{I}_{\text{ref}}$ image condition for image-to-video (I2V) generation.
$V[x:y]$ indexing operation; samples frames between frame x and (y-1).
$\text{Temporal}(w)$ temporally adjacent $w$ frames for condition.
$\text{Spatial}(T)$ spatially adjacent $T$ frames for condition.
$\tilde{V}_{\text{in}}$ conditioning frames for our framework, includes both $\text{Temporal}(w)$ and $\text{Spatial}(T)$.
$\mathcal{D}(\cdot)$ dynamic SLAM frameworks.
$\Pi(\cdot)$ perspective projection operator.
$\mathbf{P}$ aggregated 3D point clouds.
$\mathbf{P}_{\text{static}}$ aggregated 3D point clouds only from static regions.
$\mathcal{M}$ 3D scene memory composed of camera extrinsics $C_{t}$ and static point clouds $\mathbf{P}_{\text{static}}$.
$M_{i}^{\text{obj}}$ object-level masks representing dynamic regions of frame $i$.
$\mathcal{E}(\cdot)$ 3D VAE.
Appendix A Additional Experimental Results
Table 5: Quantitative evaluation of spatially consistent long video generation on DAVIS (perazzi2016benchmark) dataset.
Methods DAVIS
[…]
CameraCtrl (he2024cameractrl) 8.64 0.19 0.66 95.45 9332.08 0.1709
   + latent interpolation 14.12 0.49 0.45 53.50 3214.71 0.2225
   + FreeLong (lu2024freelong) 13.46 0.44 0.48 55.84 3314.31 0.2147
FloVD (jin2025flovd) 10.77 0.42 0.55 76.08 6131.95 0.2242
Ours 17.28 0.60 0.35 37.28 1583.77 0.1794
More Qualitative Results.
We provide additional qualitative results of our method to further demonstrate its ability to generate high-quality outputs, as shown in 9, 10, and $1$.
Long-Video Generation Results.
Although our primary interest is building a framework capable of generating a spatially consistent next-video chunk given an arbitrary video as context, by iteratively applying our method, one of the potential applications of our framework is generating videos of arbitrary length.
To generate longer videos, we extend our method using iterative autoregressive generation, and conduct quantitative evaluations on long-video generation using the DAVIS dataset (perazzi2016benchmark), where we evaluate video generation quality (PSNR, SSIM, LPIPS) and camera-controllability (RMSE, MSE, ATE), and compare with previous I2V baselines as reported in Tab. 5. However, as there are no publicly available video-to-future-video camera-controllable generation methods for our baselines, for a fairer comparison, we adopt three different strategies to generate long videos with our baseline methods. Specifically, we adopt 1) iterative autoregressive generation, 2) latent interpolation with autoregressive generation, and 3) applying training-free long video generation techniques.
For iterative autoregressive generation, we simply provide the last frame of the previously generated video as the input condition when generating the next video. For latent interpolation, we increase the number of input latents through interpolation, which can directly increase the number of generated frames at inference. After applying latent interpolation, we similarly adopt autoregressive generation to generate videos with the desired number of frames. By generating more frames at once, this reduces the number of iterations of autoregressive generation, which can increase overall video quality by reducing error accumulation. Finally, we also adopt FreeLong (lu2024freelong), a training-free approach for long video generation, which introduces frequency blending of latents specifically in Stable Video Diffusion (blattmann2023stable). As we compare with the FloVD (jin2025flovd) model fine-tuned from CogVideoX (yang2024cogvideox), which is a transformer-based model, we apply FreeLong only for CameraCtrl (he2024cameractrl). Due to latent interpolation significantly increasing the computation in CogVideoX, we also apply the latent interpolation technique for CameraCtrl only. The qualitative results of our generated long videos can be found in the supplementary videos or in Fig. 12.
As shown in the results, our method shows its potential for long video generation, significantly outperforming previous baselines. However, as is common in auto-regressive generation, extending our framework to extremely long durations introduces the challenge of temporal error accumulation. While our method provides the critical spatial consistency required for long videos, solving long-term drift (error accumulation) typically requires dedicated refinement modules and techniques. We believe that further extending our framework to mitigate the aforementioned issues is a very important and interesting future direction to explore.
Table 6: Dynamic mask ablation study.
Methods DynPose-100K
[…]
(a) w/o dynamic mask 11.9963 0.2898 0.3748 0.104002 3.4142 8.5920 10.5049
(b) (a) + L1 difference mask 12.1083 0.3248 0.3690 0.155589 2.9390 7.7819 9.1524
(c) (b) + SAM2 propagation 13.5957 0.4036 0.3548 0.157610 2.7188 7.5402 8.9696
(d) (c) + point BW tracking 13.7311 0.4112 0.3454 0.122859 2.6103 7.4858 8.9181
Ablation Study on Dynamic Masking Pipeline.
We ablate each stage of our dynamic masking pipeline presented in Sec. 3.3. Stage (a) performs training without any dynamic mask, causing moving objects to interfere with static-scene learning. Stage (b) introduces a pixel-level L1 flow-difference mask, as described in Eq.  6, which detects motion per frame but often fails to fully capture dynamic objects due to noisy or fragmented flow estimation. Stage (c) improves object-level consistency by using SAM2 propagation from points sampled in the first frame; however, this stage cannot detect new or displaced dynamic objects appearing in later frames. Stage (d) further adds point tracking across the sequence, enabling backward propagation into the first frame before SAM2 segmentation, which allows the mask to capture all dynamic objects throughout the entire clip. The quantitative comparison for each variant is summarized in Tab. 6, where we evaluate the video generation quality (PSNR, SSIM, LPIPS), scene-consistency (Met3r), and camera-controllability (mRotErr, mTransErr, mCamMC) in a subset of the DynPose-100K dataset. The results effectively verify the need for each stage of our proposed dynamic mask generation pipeline.
Table 7: Ablation study on the size of temporal window $w$. We conduct an ablation study on camera controllability with respect to the number of conditioned images $w$, and evaluate motion coherence by measuring VBench (huang2024vbench++)’s Temporal Flicker and Motion Smooth metrics. VBench metrics are normalized.
Methods DynPose-100K
[…]
Ours ($w=1$) 2.3898 7.7819 8.9785 0.8379 0.9253
Ours ($w=5$) 2.3837 7.5512 8.6233 0.8508 0.9262
Ours ($w=9$) 2.3772 7.4174 8.6352 0.8561 0.9335
Table 8: Ablation study on the number of projected images $n$. We conduct an ablation study on scene consistency and camera controllability with respect to the number of spatially adjacent frames $n$ retrieved for spatial conditioning on the DynPose-100K (rockwell2025dynamic) dataset.
Methods DynPose-100K
[…]
Ours ($n=0$) 11.9555 0.3370 0.4512 3.4142 8.5920 10.5049
Ours ($n=4$) 13.0382 0.3733 0.3758 2.3739 7.4278 8.6488
Ours ($n=7$) 13.0468 0.3666 0.3812 2.3772 7.4174 8.6352
Table 9: Using various methods for generating 3D memory.
Methods DynPose-100K
[…]
Ours w/ MegaSAM (li2024megasam) 13.0468 0.3666 0.3812
Ours w/ DepthAnything v3 (depthanything3) 13.4534 0.3980 0.3637
Ablation study on the number of temporal condition images.
As CogVideoX generates 49 frames simultaneously, the temporal window size $w$ can be selected within the range $1<w<49$, meaning (49w49-w) frames are newly generated. In our framework, setting the window size to $w=9$ (providing the last few frames) is one of the key contributions and design choices to ensure sufficient temporal context to maintain motion continuity and coherency with the input video, while still generating a sufficient number of frames (40). To better validate this choice, we conduct experiments with varying window sizes ($w=1$ and $w=5$). As summarized in Tab. 7, the results demonstrate that while the window size has a negligible effect on camera controllability metrics (mRotErr, mTransErr, mCamMC), providing a larger context window ($w=9$) yields superior motion smoothness and temporal coherence, as evidenced by improved VBench metrics (Temporal Flicker and Motion Smoothness), which is aligned with our intentions.
Ablation study on the number of projected images.
We further evaluate the impact of using different numbers of projected views as spatial conditions. To do this, we compare our method against a baseline setting ($n=0$), where $n$ denotes the number of projected frames. In this baseline, the model is conditioned solely on the temporal frames without any projected spatial views. The results, presented in Tab. 8, show that without the incorporation of our spatial prompts, the baseline framework fails to maintain scene consistency (resulting in significantly lower PSNR and SSIM scores) and struggles to adhere to the target camera trajectory. In contrast, the variants utilizing projected images ($n=4$ and $n=7$) significantly outperform the $n=0$ baseline across all metrics. This effectively verifies the critical importance of our spatial prompts for ensuring both high-fidelity scene consistency and precise camera control.
DepthAnything v3 for 3D memory.
One of the advantages of our framework is that we do not have any special architecture designs tailored to MegaSAM and can always replace MegaSAM with a more robust and powerful model to resolve the current limitations. Here, we show an experiment where we replace MegaSAM with the recently released DepthAnything v3 (depthanything3) for inference. The results, summarized in Tab. 9, show improvements across metrics, together with reduced inference time. This confirms that our method directly benefits from stronger priors and highlights a key advantage of our design: the components can be swapped at inference time without changing the overall pipeline, allowing users to flexibly choose different pretrained models.
Table 10: Inference time comparison of different methods.
Method SLAM-Processing
Dynamic Masking
& Depth Warping
Video Generation Inference Time
CameraCtrl (he2024cameractrl) 1 min 38 sec 1 min 38 sec
MotionCtrl (wang2024motionctrl) 2 min 9.5 sec 2 min 9.5 sec
FloVD  (jin2025flovd) 8 min 5.32 sec 8 min 5.32 sec
Ours – MegaSAM (li2024megasam) 4 min 18.813 sec 58.88 sec 4 min 3.62 sec 9 min 21.31 sec
Ours – DepthAnything v3 (depthanything3) 10.031 sec 58.88 sec 4 min 3.62 sec 5 min 12.53 sec
Inference Time.
Tab. 10 presents the end-to-end inference latency of our method in comparison to existing camera-controllable video diffusion models. Approaches such as CameraCtrl (he2024cameractrl), MotionCtrl (wang2024motionctrl), and FloVD (jin2025flovd) rely exclusively on diffusion-based synthesis, where inference time is primarily determined by the denoising process. In contrast, our pipeline incorporates additional stages for SLAM-based 3D reconstruction and dynamic masking. With MegaSAM (li2024megasam) employed for dynamic SLAM, the preprocessing stage requires approximately 4 minutes, resulting in a total inference time of over 9 minutes. However, replacing MegaSAM with a more lightweight and advanced method, such as DepthAnything v3 (depthanything3), significantly reduces the SLAM processing time to roughly 10 seconds. This simple replacement reduces the total latency to approximately 5 minutes, where the overall runtime is dominated by the diffusion-based video generation itself. Consequently, our framework achieves inference speeds comparable to CameraCtrl and MotionCtrl, which leverage a more lightweight video generation backbone, demonstrating that the proposed method introduces negligible computational overhead while enabling spatially consistent and camera-controllable generation.
Appendix B Training Dataset Curation Pipeline.
Our training data is curated from two primary sources: RealEstate10K (zhou2018stereo), which consists of static indoor scenes, and OpenVid-1M (nan2024openvid), which features diverse dynamic content. For static videos from RealEstate10K, we directly extract 3D scene geometry without applying dynamic masking, as these scenes contain negligible motion. In contrast, for dynamic videos from OpenVid-1M, we conduct extensive data filtering and preprocessing: we remove game-like or low-resolution videos, exclude samples with excessive camera motion, and apply the dynamic masking pipeline described in Section 3.3 to separate static and dynamic regions. We further select long video sequences with lengths of at least $L{\geq}100$ frames to ensure sufficient input video length. To construct the 3D scene memory, we employ VGGT (wang2025vggt) for static scenes and MegaSAM (li2024megasam) for dynamic scenes. After filtering and processing, our final dataset comprises approximately 30K and 20K high-quality long videos from RealEstate10K and OpenVid-1M, respectively.
Appendix C Discussion and comparison with related works
In this section, we discuss the differences between our proposed framework and recent relevant works, focusing on 3D memory integration, camera-controllable generation, and dynamic object handling.
Explicit 3D Memory and World Models
Several recent works have explored utilizing 3D memory for video generation. Persistent embodied world model (zhou2025learning3dpersistentembodied) introduces 3D memory but relies on implicit memory representations and targets static scenes. In contrast, our work leverages explicit 3D memory to achieve scene-consistent generation specifically for real-world dynamic videos. Similarly, WorldMem (xiao2025worldmemlongtermconsistentworld) builds upon the Oasis (oasis2024) architecture and is trained primarily on the Minecraft dataset. It addresses static scenes (RealEstate10K) or synthetic domains and requires key frames as conditions by increasing the sequence size. Our approach differs by leveraging the learned priors of DiTs to generate scene-consistent, camera-controllable videos in complex real-world dynamic scenarios. Furthermore, by projecting constructed static-only 3D point clouds to the target trajectory, we condition on long videos efficiently without increasing computational costs.
We also differentiate our work with SPMem (wu2025video), a concurrent work conceptually similar to ours. However, our method distinguishes itself in two key aspects. First, regarding dynamic object handling, while SPMem computes static components via naïve TSDF fusion, we introduce a concise and accurate dynamic mask generation pipeline to effectively remove dynamic regions. Second, in terms of efficiency, SPMem employs an additional ControlNet-style Diffusion-as-Shader (gu2025das) architecture, necessitating architecture search, such as the number of sufficient layers for the ControlNet-style block for different diffusion backbones, and the introduction of the additional model leads to extra computational resources for training. Conversely, we achieve scene-consistent generation without any architectural changes by injecting spatial and temporal conditioning into zero-padding slots, enabling efficient training and inference. While a direct comparison in performance and efficiency would better highlight these differences, it is currently not possible due to the unavailability of public code.
Camera-Controllable Video Generation
Works such as CameraCtrl2 (he2025cameractrl) and APT2 (lin2025autoregressive) explore camera conditioning using Plücker embeddings and autoregressive generation. While effective for short sequences, these methods typically model consistent video generation by taking only previous images as conditions without explicitly modeling 3D scene structure or considering spatial adjacency. Consequently, they fail to incorporate all previous frames due to computational constraints.
In contrast, our approach constructs an explicit 3D static memory using point clouds, enabling precise pixel-wise spatial alignment across views. Instead of relying on autoregressive propagation or feature-level Plücker conditioning, our 3D memory is directly integrated into the diffusion transformer via 3D-warped point clouds, providing scene-consistent constraints that do not accumulate drift over time. Additionally, our design explicitly handles dynamic objects by separating static and dynamic components, a challenge not addressed by these prior methods.
Dynamic Object Removal and Inpainting
Regarding the removal of dynamic objects, our spatial conditioning shares surface-level similarities with video inpainting techniques such as FGVC (gao2020flow). However, a fundamental difference lies in the role of the condition, as shown in Fig.  8. In standard inpainting, the model operates under hard constraints to fill masked regions. In our framework, spatial prompts serve as a soft spatial guide. This allows the diffusion model to fill holes and refine details while respecting the geometric layout where reliable projections exist. Crucially, this flexibility ensures the model can generate dynamic details on top of the condition, blending geometric consistency with realistic temporal dynamics, rather than merely filling static holes.
Figure 8: Visualization of the 3D scene memory and the generated results.
Appendix D Limitation
Our framework, 3DScenePrompt, enables scene-consistent, camera-controllable video generation and demonstrates improved performance over existing methods. Nonetheless, it has a few limitations. To obtain point clouds, we rely on MegaSAM (li2024megasam) to process the input videos, which introduces additional computational overhead and increases inference time. This overhead could be reduced by adopting a more advanced module, e.g., DepthAnything3 (depthanything3). In addition, the quality of the dynamic masks directly affects the fidelity of the generated videos. Thanks to our general and modular design, replacing the current component with a more advanced dynamic masking strategy that benefits from stronger models could further improve overall performance.
Figure 9: More qualitative results.
Figure 10: More qualitative results.
Figure 11: More qualitative results.
Figure 12: Long-video generation results.
Appendix E Reproducibility Statement
As mentioned in Section 3.1, our model builds upon the open-sourced CogVideoX-I2V-5B (yang2024cogvideox) model, where each of the processes for dynamic masking is also detailedly explained. We will also make all the codes publicly available.
Appendix F Use of Large Language Models
In accordance with the ICLR 2026 submission policy, we disclose that we used Large Language Models to assist in grammar correction for the writing in this manuscript.